OpenStudy (the_puppet1):

What is the solution to the equation -3 + 5x = 42? A) x = 7 B) x = 8 C) x = 9 D) x = 10

RhondaSommer (rhondasommer):

hello! I can help you :D

OpenStudy (the_puppet1):

Ok, what is it?

RhondaSommer (rhondasommer):

do you know what the first step to solve this is? if so, tell me, and write what the equation would be after that step.

OpenStudy (serenity2001):

It's c -3 + 5x = 42 +3 +3 5x = 45 /5 /5 x = 9

RhondaSommer (rhondasommer):

@serenity2001 you should know by now that direct answers (even with explination) are never allowed. Often times the person who asked the question wont even read all that hard work you took to explain it. Try quiding them through the question next time.